OpenStudy (kitten_is_back):

Add. –4.81 + (–3.64) A. –8.45 B. –1.17 C. 1.17 D. 8.45

OpenStudy (anonymous):

I can help!

OpenStudy (anonymous):

the answer id D.

OpenStudy (mathstudent55):

To add two positive numbers or two negative numbers do this: Take the absolute values of the numbers. That means if the numbers are positive, keep them positive. If the numbers are negative, make them positive by ignoring the negative signs. Add the absolute values. If the two numbers were positive, the sum is positive. If the two numbers were negative, the sum is negative.

OpenStudy (anonymous):

-8.45

OpenStudy (anonymous):

use the website www.mathway.com

OpenStudy (anonymous):

https.mathway.com

OpenStudy (like1234):

-7.82

OpenStudy (anonymous):

ok do your problem right and subtract -4.81 +(-3.64) now you do that and you get your answer and -4.81 + (-3.64) = - 8.45 and @zachary_browning please don't give direct answers cause maybe the asker needs help and not just the direct answer :)

OpenStudy (mathstudent55):

Examples: Add 5 and 7. The absolute values of 5 and 7 are 5 and 7. Add 5 + 7 = 12 Since the numbers were positive, the answer is positive, so 5 + 7 = 12 Add -6 and -8 Take the absolute values of -6 and -8. They are 6 and 8. Add 6 and 8: 6 + 8 = 14 Since the numbers were negative, the answer is negative, so -6 + (-8) = -14

OpenStudy (anonymous):

sorry I mean add

OpenStudy (anonymous):

ok ty for ur suggestiions

OpenStudy (anonymous):

your welcome @zachary_browning

OpenStudy (anonymous):

Thanks for the medal!

OpenStudy (mathstudent55):

In your case, start with the operation: –4.81 + (–3.64) You are adding two negative numbers. Take the absolute value of both numbers and add them: 4.81 + 3.64 = 8.45 Since both numbers were originally negative, the answer is negative –4.81 + (–3.64) = -8.45

OpenStudy (mathstudent55):

For this type of question, don't use a website or a calculator at first. Instead, learn how to do it. You can use a calculator afterwards to check your answer.